A tube of 20 cm length filled with a solution of 15 gm of cane
sugar in 100 cc of water is placed in the path of a polarized light. Find the angle of rotation of the
plane of polarization if the specific rotation of cane sugar is 65°.
1
Expert's answer
2020-04-30T10:39:12-0400
"[\alpha]=\frac{\alpha}{cl}"

"65=\frac{\alpha}{(2)(0.15)}"

"\alpha=19.5\degree"